Given the following cost and activity observations for Smithson Company's utilities, use the high-low method to calculate Smithson's fixed costs per month. Round your final answer to the nearest dollar. Do not round interim calculations.

Cost

Machine Hours

January $26,200 9,500
February 35,900 17,500
March 28,400 12,300
April 30,400 15,100

a.$24,958

b.$14,681

c.$11,745

d.$46,979
Answer #1

High cost – Low cost / High hours – Low Hours

=35900-26200/17500-9500 = 9700/8000 =1.2125

Total Cost = X+(1.2125*17500)

35900 = x+21218.75

X= 35900-21218.75

=14681.25

The answer is b.14681